OpenStudy (anonymous):

how do i graph a linear equation like 2x+y=4 6x+7y=12

OpenStudy (anonymous):

that is 2 equations, the easiest way for me is to solve for either x or y, usually y and plug in some values for x and find the value for y then plot points then connect the points with a line

OpenStudy (anonymous):

looking at the slope might also help

OpenStudy (anonymous):

do you wanna try it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok solve the top equation for y

OpenStudy (anonymous):

how

OpenStudy (anonymous):

get y by itself on one side of the equation

OpenStudy (anonymous):

maybe another way will be easier

OpenStudy (anonymous):

what value can you plug in for x and what value can you plug in for y that will make the expression 2x + y = 4 true, ill give you one, x = 1, y = 2, 2(1) + 2 = 4, 2 + 2 = 4, that is true so I can plot a point at (1,2)